Fermi National Laboratory


Accelerator Update

Wednesday October 29 through Friday October 31
All shifts began in shutdown.

On Wednesday, water technicians repaired the Booster LCW leak during the morning. Experts established circulating beam in the Booster and later in the evening conducted beam studies.

On Thursday, Operations ran the MiniBooNE power supplies throughout the midnight shift and then turned them off for day shift. Operators on the evening shift tried to run the MiniBooNE horn, but problems cut the test short; experts will investigate this further on Friday.

On Friday, Operation again exercised MiniBooNE power supplies throughout the midnight shift. Operators lowpotted the Main Injector buss.

TEVATRON STATUS
Sector E17 conning tower connection fix finished today (10/31/03)
Sector A0 Antiproton kicker girder move completed today (10/31/03)
Finish the HLS water level system in B sector today (10/31/03)
Install hasps and locks on survey vents by the end of today (10/31/03)
Complete the BPM work and cable installation by the end of week (10/31/03)
Replace flying wire at E17 by the end of today (10/31/03)
Complete the power connection of the A0 sextupole and its polarity check and controls test today (10/31/03). Here is a picture of the newly installed sextupole.

TEVATRON Future Schedule
For the week ending November 7th

  • Final survey for the sector A-48 collimator
  • Move the sector D0 low beta quad girder
  • Conduct a final laser tracker survey check of the sector F0 Lambertson
  • Begin checking smart bolt cold lifts after the TeV cryo system cools the magnets down to liquid helium temperature, which should happen near the end of the week.

For the week ending November 14th

  • Check the Collider Detector Facility (CDF) low beta as founds
  • Install poly beads at CDF sometime after November 11th
  • Complete all fire stop fills

RECYCLER STATUS

  • 24 or 28 ring bakes have been completed. The project is 86% complete.
  • The alignment of Recycler components is 80% complete.
  • The Rework of the 620 region is 80% complete.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/31/2003   email Fermilab